### v1

# HiFo-Prompt: Prompting with Hindsight and Foresight for LLM-based Automatic Heuristic Design

Abstract
LLM-based Automatic Heuristic Design (AHD) within Evolutionary Computation (EC) frameworks has shown promising results. However, its effectiveness is hindered by the use of static operators and the lack of knowledge accumulation mechanisms. We introduce HiFo-Prompt, a framework that guides LLMs with two synergistic prompting strategies: Foresight and Hindsight. Foresight-based prompts adaptively steer the search based on population dynamics, managing the exploration-exploitation trade-off. In addition, hindsight-based prompts mimic human expertise by distilling successful heuristics from past generations into fundamental, reusable design principles. This dual mechanism transforms transient discoveries into a persistent knowledge base, enabling the LLM to learn from its own experience. Empirical results demonstrate that HiFo-Prompt significantly outperforms state-of-the-art LLM-based AHD methods, generating higher-quality heuristics while achieving substantially faster convergence and superior query efficiency.
1 Introduction
Combinatorial Optimization (CO) is a significant branch of mathematical optimization (dreo2006metaheuristics, ), focusing on identifying solutions that maximize or minimize an objective function over a discrete feasible space.
Over the past decades, numerous efficient heuristic algorithms (BlumRoli2003, ; ColorniEtAl1996, ) have emerged in the field of optimization to solve these problems.
However, manual design of heuristic algorithms remains a task that heavily relies on domain expertise, requiring a deep understanding of problem structures and extensive experiential knowledge (CamachoVillalonEtAl2023, ).
In recent years, the rise of Large Language Models (LLMs) has brought new opportunities for the automated design of heuristic algorithms (VaswaniEtAl2017, ; WangChen2023, ; LiuYaoEtAl2024, ). A number of LLM-based Automated Heuristic Design (LLM-based AHD) methods have been proposed, many of which apply the Evolutionary Computation (EC) frameworks. Pioneering works such as FunSearch (romera2024, ) and EoH (liu_icml2024, ) have demonstrated the effectiveness of the evolutionary computation paradigm. FunSearch uses LLMs to evolve programs, searching for high-quality heuristic functions in a predefined heuristic program space. EoH introduces the co-evolution of thought and code, enabling LLMs to iterate between the thought process and code generation, thereby simulating the heuristic design process employed by human experts.
Following prior works, some LLM-based AHD approaches (vanSteinBack2025, ; ZhangLiuEtAl2024, ) focus on designing diverse guidance methods to enhance the search process. For example, ReEvo (ye2024, ) proposes a reflection-based framework to simulate language gradient guidance for directing the search process. HSEvo (Dat2025HSEvo, ) integrates diversity metrics with harmony search to balance population diversity and convergence better. However, these methods remain inadequate in terms of knowledge management wu2024 , as they do not incorporate a systematic mechanism for explicitly distilling and reusing algorithmic principles proven effective during evolution. Consequently, valuable insights generated by LLMs are often transient, being either discarded with the elimination of parent candidates or diluted over successive iterations.
Meanwhile, some LLM-based AHD approaches seek breakthroughs by refining the structure of the search process. MCTS-AHD (zheng2025, ) uses Monte Carlo Tree Search to manage heuristics, retaining temporarily underperforming but promising individuals to enable broader exploration. However, such methods typically embed regulation mechanisms passively within fixed algorithmic workflows, with knowledge utilization confined to local reasoning on individual population members. They lack an independent system capable of monitoring the global population’s state and proactively switching generation modes upon detecting issues.
In this paper, to address the limitations of the above approaches, we propose a novel framework named Hindsight-Foresight Prompt (HiFo-Prompt). At the individual level, we decouple thought from code and independently extract each individual’s thought, referred to as an insight. Then, these insights are added to an Insight Pool, a dynamic repository that accumulates and updates distilled knowledge during evolution. This process improves the framework’s Hindsight capability. At the population level, we introduce the Evolutionary Navigator, a control function that monitors the global state of the population after each iteration. Based on this state information, it adaptively determines the exploration–exploitation mode for the next iteration, strengthening the Foresight ability of the framework.
In summary, the contributions of our approach are as follows:
  • We introduce HiFo-Prompt, a novel framework consisting of Hindsight and Foresight modules. It dynamically generates prompts for LLMs by decoupling thoughts from code, thereby enabling independent updates and evaluations. This mechanism leads to a significant reduction in both training time for heuristics and evaluation costs for LLMs.
  • To improve the Hindsight and Foresight abilities of our method, we introduce the Insight Pool and Evolutionary Navigator, respectively. The Insight Pool accumulates knowledge from high-performing codes through iterative updates. The Evolutionary Navigator controls population states by monitoring evolution and balancing exploration-exploitation dynamics.
  • We evaluated the heuristics designed by HiFo-Prompt on complex optimization tasks, comparing them against advanced handcrafted heuristics and existing AHD approaches. Our results achieve state-of-the-art performance in the AHD domain, with substantial improvements over prior AHD methods, particularly excelling in the Traveling Salesman Problem (TSP) and Flow Shop Scheduling Problem (FSSP).
2 Related Work
2.1 LLM-based Automatic Heuristic Design
AHD (PotterCulleyDarlingtonChawdhry2003, ) has traditionally relied on Hyper-Heuristics (HH) for CO (drake2020, ; burke2009, ; DokerogluKucukyilmazTalbi2024, ). However, HH’s dependence on generating or selecting from handcrafted components (burke2009, ; burke2018, ) inherently constrains algorithmic novelty and adaptability (li2024, ). While the advent of LLMs opened new frontiers, initial work showed that naive prompting often fails to produce competitive heuristics for complex problems (yang2024, ; kristiadi2024, ; guo2024, ).
A more promising paradigm integrates LLM with EC, a synergy pioneered by Funsearch (romera2024, ). In this LLM with EC approach, the LLM acts as a sophisticated genetic operator (i.e., crossover and mutation). This method has fueled successful frameworks like EoH (liu_icml2024, ) and its successors, which have been rapidly extended to domains like code generation and heuristic design (ye2024, ; zheng2025, ; liu2023, ; zhang2025, ).
2.2 Knowledge Management in Generative Search
Leveraging historical information is critical for efficient search. In classical EC, Cultural Algorithms (maher2021comprehensive, ; reynolds2000cultural, ) employ a "Belief Space" to store and evolve collective knowledge, guiding population evolution in a structured manner (reynolds2000cultural, ). In the era of LLMs, this concept has been mirrored by "reflection" mechanisms, where models refine their outputs based on verbal feedback (shinn2023reflexion, ; bo2024reflectivemultiagent, ). However, this self-reflection often produces unstructured and volatile feedback that is difficult to reuse systematically across generations (zhang2024selfcontrast, ). Furthermore, it can be biased by the model’s pre-trained knowledge rather than objective empirical evidence from the ongoing search (zhang2024selfcontrast, ).
2.3 Adaptive Control in Evolutionary Computation
Dynamically adapting strategies based on the search state is a long-standing goal in EC. Traditional methods include fine-grained adjustments such as Adaptive Operator Selection (AOS) (fialho2010adaptive, ; Tian2023Deep, ) and parameter control (eiben2015, ; aleti2016systematic, ). Recently, researchers have begun to use LLMs’ reasoning capabilities for adaptive interventions, such as generating diverse individuals to escape local optima (yu2024deep, ; veerapen2012exploration, ). This marks a conceptual shift from reactive, low-level parameter tuning to proactive, high-level strategic reasoning, allowing the search to adapt based on a semantic understanding of the population’s state (Eiben1999, ; Karafotias2015, ; Papa2021Applications, ; deLacerda2021systematic, ). However, the efficacy of these interventions often relies on predefined triggers and static prompts, which may not be optimal across all stages of the evolutionary process (Zhang2012Survey, ; Kramer2010Evolutionary, ).
3 Methodology
In this section, we introduce our proposed HiFo-Prompt, a novel framework that equips the LLM-driven evolutionary process with mechanisms for learning and adaptation (shown in Figure 1). HiFo-Prompt integrates two synergistic components: a Foresight module for real-time adaptive control, which monitors evolutionary dynamics to steer the search strategy, and a Hindsight module for long-term knowledge accumulation, which manages a self-evolving repository of successful design principles that we term insights. By uniting foresight-driven strategy with hindsight-informed knowledge, our framework transforms the generative process into a robust, self-regulating system.
Figure 1: The framework of HiFo-Prompt. Left: Prompt Construction builds upon a foundational prompt, augmenting it with strategic Foresight and empirical Hindsight to form the final composite prompt. Right: Knowledge Evolution forms a learning loop where elite heuristics are distilled into new insights to continuously enrich the Hindsight knowledge base.
3.1 Guided Prompt Synthesis for AHD
Our HiFo-Prompt framework applies a Guided Prompt Synthesis mechanism that constructs each prompt as a context-aware composite instruction. This mechanism integrates three interlocking modules: Foundational prompt strategies, hint knowledge, and a foresight-driven strategy. The resulting composite prompt provides precise, multifaceted guidance to the LLM. A complete example of prompt generation for TSP with step-by-step construction is provided in Figure 2.
Figure 2: Dynamic prompt generation process of HiFo-Prompt.
Foundational prompt strategies.
Our framework’s generative foundation is a set of Foundational Prompt Strategies, which function as the LLM-equivalent of genetic operators. We first generate heuristics from scratch using the initial prompt strategy I1, then evolve them with five foundational prompts adapted from EoH liu_icml2024 . These prompts are organized into two primary strategies: 1) Reorganization Strategies, which include E1, synthesizing a new algorithm with a novel structure from multiple parents, and E2, abstracting shared core ideas to generate conceptually distinct variants; and 2) Mutation Strategies, which encompass M1, making structural modifications for functionally equivalent variants; M2, tuning critical parameters; and M3, simplifying components prone to overfitting. While this curated set provides the raw generative capability, its effectiveness depends on contextual guidance. This is the role of the Hindsight and Foresight modules, which inject insights and a design directive into the prompts to align each action with the search’s current needs.
Hindsight Module.
This module incorporates proven knowledge in the form of insights, which are abstract and generalizable design principles distilled from successful heuristics. These insights are managed in a dynamic Insight Pool, where each is assigned and continuously updated with a credibility score based on its empirical performance. Before generation, high-scoring insights are retrieved and embedded into the prompt. They serve as validated priors to steer the LLM toward promising designs. While effective for historical guidance, this module cannot address real-time evolutionary needs.
Foresight Module.
The Foresight module implements real-time strategic control, orchestrated by its core component, the Evolutionary Navigator. The Navigator continuously monitors macroscopic evolutionary indicators, such as performance stagnation and population diversity, to assess the state of the search. Based on this analysis, it selects the governing evolutionary regime for the subsequent generation. This regime dictates the overall strategic focus by choosing one of three explicit modes (Crepinsek2013Exploration, ): 1) Explore, to foster novelty when diversity is low or progress has stalled; 2) Exploit, to refine high-performing solutions when progress is consistent; 3) Balance, to maintain a synergistic application of all operators. This high-level directive shapes the final prompt, ensuring that the generative process of the LLM aligns with the immediate strategic needs of the evolution.
3.2 Hindsight: Mechanisms of the Self-Evolving Insight Pool
While foundational prompts provide the generative actions, the Hindsight module steers evolution with empirically validated knowledge managed within the Insight Pool. Analogous to the Belief Space in Cultural Algorithms (coello2010efficient, ), our work introduces a novel instantiation for LLM-based AHD, distinguished by two core features: LLM-driven insight extraction and a utility-based credit assignment tailored for prompt-based generation.
To mitigate the cold-start problem, we initialize the Insight Pool with seed insights from established heuristic design literature. The core objective of the framework then becomes the autonomous discovery and refinement of new, problem-specific insights. This is achieved through a continuous lifecycle that systematically transforms transient evolutionary successes into explicit, reusable knowledge assets, governed by three integrated phases: insight extraction, utility-driven application, and adaptive pruning.
Insight Extraction and Admission.
The lifecycle begins by expanding the knowledge base. At the end of each generation, we prompt an LLM to distill generalizable design principles from the elite individuals of the population (see Appendix G.1). To preserve informational diversity, these candidate insights are admitted only if their Jaccard similarity to all existing pool members falls below a novelty threshold, thus preventing redundancy. Admitted insights become active candidates for guiding future generations.
Insight Retrieval and Credit Assignment.
To guide heuristic generation, we employ a utility-based retrieval mechanism. Instead of random sampling, it selects the top-$s$ insights with the highest adaptive utility score $U(k_{i},t)$. Subsequently, we use a credit assignment (Whitacre2006Credit, ) to update these utility scores, rewarding insights that lead to high-performing heuristics. The utility function is formulated to balance exploitation and exploration:
$U(k_{i},t)=\underbrace{E_{i}(t)}_{\text{Effectiveness}}-\underbrace{w_{u}\log(N_{i}(t)+1)}_{\text{Usage Penalty}}+\underbrace{B_{r}(t,\tau_{i})}_{\text{Recency Bonus}}$ (1)
where $E_{i}(t)$ is the learned effectiveness, the penalty term (weighted by $w_{u}$) discourages overuse, and the recency bonus $B_{r}$ promotes strategic coherence by rewarding insights used in recent successful generations.
The effectiveness score $E_{i}(t)$ is updated via credit assignment. This process first converts the raw fitness $g(h_{\text{new}})$ of an offspring into a normalized, problem-agnostic score $\tilde{\rho}$, scaling its performance relative to the current population’s best and worst solutions:
$\tilde{\rho}=\frac{g(h_{\text{worst}})-g(h_{\text{new}})}{g(h_{\text{worst}})-g(h_{\text{best}})}$ (2)
We posit that an offspring’s evolutionary contribution is non-linear. Therefore, we map $\tilde{\rho}$ to a final credit signal, $g_{\text{eff}}$, using a tiered, piecewise function. This deliberate design creates distinct reward regimes for qualitatively different outcomes:
$g_{\text{eff}}=\begin{cases}0.8+0.2\cdot\tilde{\rho}&\text{if }g(h_{\text{new}})\leq g(h_{\text{best}})\\ 0.2+0.6\cdot\tilde{\rho}&\text{if }g(h_{\text{best}})<g(h_{\text{new}})\leq g(h_{\text{avg}})\\ -0.3+0.5\cdot\tilde{\rho}&\text{if }g(h_{\text{new}})>g(h_{\text{avg}})\end{cases}$ (3)
This structure provides strong positive signals for paradigm shifts, moderate rewards for incremental progress, and penalties for below-average performance. Finally, for each insight $j$ in the contributing set $K_{c}$, its effectiveness score is updated through an Exponential Moving Average (EMA):
$E_{j}(t+1)=(1-\alpha)\cdot E_{j}(t)+\alpha\cdot g_{\text{eff}}$ (4)
where the learning rate $\alpha\in[0,1]$ is set to a small value to smooth the stochastic credit signals from individual evaluations. This allows an insight’s robust, long-term utility to emerge statistically, preventing its score from being skewed by outlier performances.
Adaptive Pruning and Pool Maintenance.
To maintain quality within its finite capacity $C_{\text{pool}}$, the pool employs an adaptive pruning mechanism triggered when $|K_{\text{pool}}|>C_{\text{pool}}$. This process removes the insight with the minimum eviction score, $\mathcal{S}_{\text{evict}}$, which balances an insight’s proven performance against the risk of its obsolescence:
$\mathcal{S}_{\text{evict}}(k_{i},t)=E_{i}(t)-R_{\text{decay}}\cdot(t-t_{\text{last\_used}}(k_{i}))$ (5)
where $E_{i}(t)$ is the current effectiveness of the insight, $t_{\text{last\_used}}(k_{i})$ is the generation of its last use, and the time decay rate $R_{\text{decay}}$ targets not only low-effectiveness insights, but also those that have become inactive. The decay rate is set conservatively to prevent the premature removal of valuable but temporarily dormant knowledge. Furthermore, a grace period grants new insights with usage counts below a threshold $T_{\text{usage}}$, eviction immunity. This ensures novel principles are given a fair opportunity to demonstrate their utility before facing pruning.
3.3 Foresight: The Evolutionary Navigator for State-Aware Guidance
While the Hindsight module grounds the search process in historically validated principles, we propose the Foresight module, an Evolutionary Navigator that provides real-time, state-aware guidance for heuristic design. Its primary role is to dynamically regulate the balance between exploration and exploitation throughout the evolutionary process. It manages the evolutionary strategy by implementing a control policy $\pi$ that maps the macroscopic evolutionary state, $S_{t}$, to a high-level strategic orientation known as the Evolutionary Regime, defined as follows:
$\theta_{t}=\pi(S_{t})$ (6)
The control policy $\pi$ employs a rule-based logic to determine $\theta_{t}$, prioritizing responses to critical evolutionary dynamics such as performance stagnation or diminishing diversity. The macroscopic state $S_{t}$ encapsulates these dynamics through key metrics, formalized as:
$\theta_{t}=\begin{cases}\theta_{\text{explore}}&\text{if }C_{\text{stag}}(t)\geq T_{\text{stag}}\text{ or }\Delta_{p}(t)<\Delta_{\text{crit}}\\ \theta_{\text{exploit}}&\text{if }C_{\text{prog}}(t)\geq T_{\text{prog}}\\ \theta_{\text{balance}}&\text{otherwise}\end{cases}$ (7)
where $C_{\text{stag}}(t)$ and $C_{\text{prog}}(t)$ are counters for consecutive generations of stagnation and progress, respectively, and $\Delta_{p}(t)$ measures population diversity. Specifically, stagnation is defined as a lack of improvement in the best-so-far fitness, while progress requires a significant fitness gain. Population diversity, $\Delta_{p}(t)$, is quantified as the standard deviation of the fitness values within the current population.
The hyperparameters $T_{\text{stag}}$, $T_{\text{prog}}$, and $\Delta_{\text{crit}}$ are thresholds that ensure the Navigator is highly responsive to the evolving search landscape. The thresholds for stagnation ($T_{\text{stag}}$) and progress ($T_{\text{prog}}$) are intentionally set to small values, a design choice motivated by the rapid convergence of our framework. Unlike traditional Evolutionary Algorithms that may evolve over many generations, HiFo-Prompt often achieves significant progress in very few iterations, necessitating a Navigator that can react swiftly to these short-term dynamics. Similarly, the critical diversity threshold, $\Delta_{\text{crit}}$, is carefully calibrated to balance the need to maintain population diversity to converge on high-quality solutions.
Once the strategic regime is determined, the Evolutionary Navigator translates this strategy into a concrete instruction for the LLM, known as a Design Directive. The nature of this directive is dictated by the active regime ($\theta_{t}$). For example, the $\theta_{\text{explore}}$ regime yields directives that prompt radical innovation, such as invent a new crossover operator from scratch. Conversely, the $\theta_{\text{exploit}}$ regime issues directives focused on incremental refinement, like slightly modify the selection pressure of the best heuristic. The $\theta_{\text{balance}}$ regime generates directives that synthesize these approaches, such as combine the strengths of the top two performing heuristics. This mechanism ensures that the generative focus of the LLM is dynamically steered to align with the strategic intent of the current evolutionary phase.
4 Experiments
Method TSP10 TSP20 TSP50 TSP100 TSP200
Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj.
LKH3 0.000% 2.846 0.000% 3.863 0.000% 5.709 0.000% 7.793 0.000% 10.727
EoH 6.480% 3.030 9.236% 4.220 11.553% 6.369 13.638% 8.856 14.209% 12.251
ReEvo 5.215% 2.994 6.156% 4.101 9.714% 6.264 11.379% 8.680 12.830% 12.103
HSEvo 5.075% 2.990 7.320% 4.146 11.185% 6.348 11.944% 8.724 13.463% 12.171
MCTS-AHD 4.829% 2.983 8.045% 4.174 10.642% 6.317 12.521% 8.769 13.510% 12.176
Ours 1.560% 2.890 2.790% 3.971 5.808% 6.041 8.158% 8.429 9.315% 11.726
Table 1: Results on TSP with step-by-step construction. Gap(%) denotes the performance gap compared to advanced heuristic algorithms. Obj. represents the value of the objective function. The best results are highlighted in bold.
Method TSP50 TSP100 TSP200
POMO 0.189% 0.496% 2.935%
LEHD 0.099% 0.131% 0.342%
EoH 0.000% 0.026% 0.453%
HSEvo 0.000% 0.030% 0.821%
Ours 0.000% 0.014% 0.397%
Table 2: Results on TSP with GLS. Comparison relative to the results of advanced heuristic on TSP50, TSP100 and TSP200. The best-performing overall and LLM-based AHD results are shown in bold.
In this section, we present the results of heuristics designed by our proposed HiFo-Prompt on different complex tasks, including Traveling Salesman Problem (TSP)(Matai2010, ), Online Bin Packing Problem (Online BPP) (Seiden2002Online, ), Flow Shop Scheduling Problem (FSSP) (Emmons2012FlowShop, ), and Bayesian Optimization (BO) (Shahriari2016Taking, ). Task definitions and details are given in Appendix B. Results on TSP, online BPP, and FSSP are presented in this section, while the results on BO are provided in Appendix C.4, where HiFo-Prompt demonstrates competitive and reliable performance.
Settings.
We use the qwen2.5-max model via the DashScope API with a temperature of 1.0 and evolve a population of 4 individuals. The Insight Pool in Hindsight module has a capacity of $C_{\text{pool}}=30$, with a Jaccard similarity threshold of $0.7$ for deduplication. Insight retrieval uses a utility function with selection count $s=3$, usage penalty weight $w_{u}=0.1$, and recency bonus $\tau_{r}=0.2$. Credit assignment is performed using EMA with a learning rate of $\alpha=0.3$. For pool maintenance, each insight’s eviction score decays at $R_{\text{decay}}=0.01$ per generation, and new insights receive a grace period of $T_{\text{usage}}=3$ applications. In Foresight module, Evolutionary Navigator’s thresholds are set to $T_{\text{stag}}=3$ for stagnation, $T_{\text{prog}}=2$ for progress, and $\Delta_{\text{crit}}=0.3$ for critical diversity. We run 8 generations for CO tasks and 4 for BO, where rapid convergence demonstrates our method’s efficiency. To ensure fairness, all LLM-based AHD baselines use the same LLM endpoint and query budget, with other settings kept as reported in their original papers.
Baselines.
To demonstrate the effectiveness of our proposed method in designing heuristics, we introduce several approaches for solving these complex optimization tasks. (1) handcrafted heuristics, e.g., LKH3(LinKernighan1973, ) for TSP, First Fit and Best Fit (romera2024, ) for Online BPP, NEH (nawaz1983heuristic, ) and NEHFF (FernandezViagas2014, ) for FSSP. (2) Neural Combinatorial Optimization (NCO) methods, e.g., POMO (kwon2020pomo, ) and LEHD (luo2023neural, ) for TSP, PFSPNet_NEH (pan2022deep, ) for FSSP. (3) LLM-based AHD methods, e.g., Funsearch (romera2024, ), EoH (liu_icml2024, ), ReEvo (ye2024, ), HSEvo (Dat2025HSEvo, ) and MCTS-AHD (zheng2025, ). Most of our test datasets follow EoH. Notably, Funsearch, ReEvo, and HSEvo require a seed function to initialize their populations, while EOH, MCTS-AHD, and our method can run without it.
Method 5k_C100 5k_C300 5k_C500
First Fit* 4.40% 4.18% 4.27%
Best Fit* 4.08% 3.83% 3.91%
Funsearch* 0.80% 1.07% 1.47%
EOH 1.07% 1.07% 1.06%
ReEvo 0.78% 4.47% 3.24%
HSEvo 0.79% 5.41% 4.04%
MCTS-AHD 0.99% 0.95% 0.95%
Ours 0.66% 0.63% 0.66%
Table 3: Results on Online BPP. The results denote the ratio of excess bins compared to the lower bound on Weibull instances. Results with * are from EoH (liu_icml2024, ). The best results are highlighted in bold.
Traveling Salesman Problem.
We design heuristics for solving the TSP using both Step-by-Step Construction (asani2023convex, ) and Guided Local Search (GLS) (alsheddy2016guided, ). Table 1 compares the results of our method with other baselines. We evaluated the performance on 100 instances at each of five sizes. To demonstrate performance on out-of-distribution instances, we also conducted experiments on the TSPLib dataset (reinelt1991tsplib, ), and the results can be found in the Appendix C.1. We further evaluated our method on 100 instances of TSP50, TSP100, and TSP200 in GLS. The results are shown in Table 2. The heuristic designed by our method achieves a significant performance improvement compared to the LLM-based AHD approach. It can even outperform or achieve comparable performance to some outstanding NCO methods (vinyals2015pointer, ) on small-scale instances.
nn20,mm10 nn50,mm10 nn100,mm10
NEH* 4.05% 3.47% 2.07%
NEHFF* 4.15% 3.62% 1.88%
PFSPNet_NEH* 4.04% 3.48% 1.72%
EOH 0.244% 0.287% 0.174%
Ours 0.145% 0.120% 0.094%
nn20,mm20 nn50,mm20 nn100,mm20
NEH* 3.06% 5.48% 3.58%
NEHFF* 2.72% 5.10% 3.73%
PFSPNet_NEH* 2.96% 5.05% 3.56%
EOH 0.179% 0.753% 0.927%
Ours 0.103% 0.600% 0.472%
Table 4: Results on FSSP. The results show the comparison of makespan relative to the baseline on Taillard instances. $n$ denotes the number of jobs and $m$ denotes the number of machines. Results with * are from EoH (liu_icml2024, ). The best results are highlighted in bold.
TSP
TSP20 TSP50 TSP100
Gap Obj. Gap Obj. Gap Obj.
w/o Insight Pool 11.07% 4.29 13.76% 6.50 16.26% 9.06
w/o Navigator 5.82% 4.09 10.31% 6.30 11.48% 8.69
w/o Insight Pool and Navigator 11.49% 4.31 14.36% 6.53 18.80% 9.26
HiFo-Prompt 2.79% 3.97 5.81% 6.04 8.16% 8.43
Online BPP
1k, 100 5k, 100 1k, 300 5k, 300 1k, 500 5k, 500
w/o Insight Pool 4.33% 1.27% 4.07% 1.22% 4.14% 1.29%
w/o Navigator 2.83% 1.26% 2.64% 1.24% 2.60% 1.24%
w/o Insight Pool and Navigator 4.53% 2.19% 4.30% 2.01% 4.26% 2.05%
HiFo-Prompt 1.99% 0.66% 2.01% 0.63% 1.95% 0.66%
Table 5: Ablations on Insight Pool and Navigator in TSP and Online BPP. The best results are highlighted in bold.
Online Bin Packing Problem.
The key function of Online BPP is a scoring function that outputs a score for each bin, based on the current item’s size and the remaining capacities of the bins. We evaluate our method on Weibull BPP instances (romera2024, ) following the EoH setting. The results with three scales are shown in Table 3. Our method not only significantly outperforms handcrafted heuristics but also surpasses LLM-based AHD approaches. More results are provided in Appendix C.2.
Flow Shop Scheduling Problem.
FSSP involves scheduling $n$ jobs on $m$ machines to minimize the makespan, where each job comprises $m$ operations processed in a fixed order. We evaluate the heuristic designed by our method on Taillard instances (taillard1993benchmarks, ). The dataset includes instances with 20 to 100 jobs ($n$) and 10 to 20 machines ($m$). Table 4 presents some of the results, with additional results provided in Appendix C.3. Our method performs well on all datasets, consistently delivering strong and reliable results across different scenarios.
Figure 3: Convergence Analysis for TSP and Online BPP.
Ablation Study.
To evaluate the contribution of each component in our method, we conducted a series of ablation studies, as shown in Table 5. We separately removed the Hindsight obtained by the insight pool and the Foresight provided by the navigator, and then removed both Hindsight and Foresight entirely. These experiments are conducted on TSP and Online BPP. The design and parameters of the experiments are aligned with those used in the main experiments.
Comparative Results.
We compared the progression of objective function values during the evolutionary process of our method and EoH on both the TSP and Online BPP. Figure 3 presents the convergence analysis curves. The figure demonstrates that our method converges more rapidly while requiring fewer individuals in the population, indicating higher efficiency. In addition, to further highlight the efficiency of our method in terms of training time and computational cost, we measured both the training duration and the number of LLM requests for the four methods. The results are presented in Figure 4.
Figure 4: Efficiency Comparison on LLM Requests and Training Time for TSP and Online BPP.
5 Conclusion
We introduced HiFo-Prompt, a novel evolutionary framework that advances LLM-driven heuristic design through a hierarchical foresight-hindsight prompting mechanism. By synergizing an Evolutionary Navigator for adaptive control with a self-evolving Insight Pool for knowledge reuse, our framework transforms the search process into a closed-loop, self-regulating system. Across diverse optimization benchmarks, HiFo-Prompt consistently outperforms state-of-the-art methods with remarkable sample efficiency, often finding superior solutions using only 200 LLM requests. This work provides not only a powerful method for heuristic automated algorithm design but also a concrete step towards agents that can learn to invent their problem-solving methodologies.
Appendix A Preliminary
A.1 Problem Formulation of Automatic Heuristic Design
Automatic Heuristic Design (AHD) [18, 62] aims to identify an optimal heuristic $h^{*}$ from a vast search space $\mathcal{H}$ for a given computational task $\mathcal{P}$. This process can be formally expressed as the following optimization problem [15]:
$h^{*}=\mathop{\arg\max}_{h\in\mathcal{H}}g(h),$ (8)
where $g(h)$ is a fitness function that maps a heuristic to a real number, estimating its quality based on its expected performance on a representative set of problem instances $D$. For a task with a minimization objective $f$ (e.g., minimizing cost or error), this performance metric is defined as:
$g(h)=\mathbb{E}_{\mathrm{ins}\in D}\left[-f(h(\mathrm{ins}))\right].$ (9)
This formulation reframes the original task as a maximization problem over the heuristic space $\mathcal{H}$, thereby enabling the search for robust heuristics that yield high-quality solutions across diverse instances.
A.2 LLM-driven Evolutionary Computation
The LLM-driven Evolutionary Computation (LLM+EC) framewor[9, 8, 63, 64] casts the AHD problem as an iterative, population-based search process. Let $P^{(t)}=\{h_{1}^{(t)},\dots,h_{M}^{(t)}\}$ be the population of $M$ heuristics at generation $t$, where each heuristic $h_{i}^{(t)}\in\mathcal{H}$ is represented by its thought and code. The transition from $P^{(t)}$ to $P^{(t+1)}$ is governed by a stochastic evolutionary kernel $\mathcal{F}$, which is parameterized by a control vector $\theta^{(t)}\in\Theta$:
$P^{(t+1)}\sim\mathcal{F}(P^{(t)}\mid\theta^{(t)}).$ (10)
Here, the control vector $\theta^{(t)}$ encapsulates contextual information that guides heuristic generation, such as prompting strategies or performance feedback. In the absence of adaptive control, $\theta^{(t)}$ can be considered constant or null, i.e., $\theta^{(t)}=\theta_{\text{const}}$ or $\theta^{(t)}=\varnothing$. The kernel $\mathcal{F}$ comprises two phases:
  1. 1.
    Generation Phase: An LLM, acting as a conditional generator $\mathcal{L}$, creates new heuristics through prompted crossover and mutation. A set of parents $h_{p}\subseteq P^{(t)}$ is selected, and their symbolic representations $\rho(h_{p})$, which include the thought-and-code, are used. A prompt function $\Pi$ constructs a conditional prompt from $\rho(h_{p})$ and $\theta^{(t)}$ (e.g., exploration or modification strategies). The LLM then generates offspring:
    $h_{c}=\mathcal{L}(\Pi(\rho(h_{p}),\theta^{(t)})),\quad O^{(t)}=\{h_{c,1},h_{c,2},\dots,h_{c,N}\},$ (11)
    where $N$ is the number of offspring (e.g., $N=\lambda M$, where $\lambda$ is the reproduction rate).
  2. 2.
    Selection Phase: The parent and offspring populations are merged into a candidate pool, $U^{(t)}=P^{(t)}\cup O^{(t)}$. A selection operator $\mathcal{S}$ then chooses the top $M$ heuristics from this pool based on the fitness metric $g$ to form the next generation:
    $P^{(t+1)}=\mathcal{S}(U^{(t)};g).$ (12)
This framework, introduced within the "heuristic evolution" paradigm, leverages LLM-driven generation to explore the heuristic space and a selection mechanism to exploit high-performing solutions.
A.3 Knowledge-Augmented Evolutionary Computation
To extend the capabilities of evolutionary algorithms beyond simple adaptive control, a prominent research direction involves incorporating an explicit knowledge component. Cultural Algorithms [28, 45, 65] provide a classic framework for this idea, decoupling the evolutionary system into two interacting spaces: a population space containing candidate solutions $P^{(t)}$ and a belief space $K_{t}$ serving as a repository of experiential knowledge [27]. These two spaces co-evolve through a dual-inheritance communication protocol.
Crucially, knowledge $k_{s}$ extracted from the belief space $K_{t}$ becomes part of the high-level control vector $\theta^{(t)}$. This knowledge then guides the generation of offspring via an influence mechanism:
$h_{c}\sim\mathcal{L}\bigl{(}\Pi(\rho(h_{p}),\theta^{(t)})\bigr{)},\quad\text{where }k_{s}\subseteq\theta^{(t)}.$ (13)
Concurrently, the successes of the population are fed back into the belief space. An acceptance function $\mathcal{A}$ identifies and extracts potentially valuable experiences, $\mathcal{A}(P^{(t)})$, from the current population. These are then used by a knowledge update function $\mathcal{U}_{K}$ to update the belief space:
$K_{t+1}=\mathcal{U}_{K}(K_{t},\mathcal{A}(P^{(t)})).$ (14)
Knowledge-augmented evolution is thus a coupled dynamical system where the population and knowledge base co-evolve interdependently. This explicit mechanism for knowledge management allows the framework to achieve a more sophisticated form of learning based on abstracted experience.
Appendix B Optimization Problem Details
B.1 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) [47, 66] is a canonical NP-hard combinatorial optimization problem. Given a set of $N$ cities and the distances between each pair, the objective is to find the shortest possible tour that visits each city exactly once before returning to the starting city [67].
Formally, let $V=\{v_{1},v_{2},\dots,v_{N}\}$ be the set of cities. We model the problem on a complete, undirected graph $G=(V,E)$, where a non-negative distance $d_{ij}$ is associated with each edge $(v_{i},v_{j})\in E$. A tour is a permutation $\pi$ of the indices $\{1,2,\dots,N\}$. The goal is to find a permutation $\pi^{*}$ that minimizes the total tour length [68]:
$\min_{\pi}\left(\sum_{i=1}^{N-1}d_{\pi_{i},\pi_{i+1}}+d_{\pi_{N},\pi_{1}}\right)$
where $v_{\pi_{i}}$ denotes the $i$-th city in the tour.
Step-by-step Construction.
Construction heuristics build a feasible solution from an empty set by making a sequence of decisions [69]. At each step, a component is added to the partial solution based on a specific greedy criterion until a complete tour is formed [70]. A fundamental example is the Nearest Neighbor (NN) heuristic. Starting from a node $v_{\pi_{1}}$, it constructs a tour by iteratively selecting the closest unvisited node. At step $t$, with a partial tour $S_{t}=(v_{\pi_{1}},\dots,v_{\pi_{t}})$ and the set of unvisited nodes $U_{t}=V\setminus\{v_{\pi_{1}},\dots,v_{\pi_{t}}\}$, the next node $v_{\pi_{t+1}}$ is chosen according to the rule [71]:
$v_{\pi_{t+1}}=\arg\min_{v_{j}\in U_{t}}d_{\pi_{t},j}$
While fast, the myopic nature of such simple rules often leads to suboptimal solutions.
Our approach, HIFO-Prompt, introduces a new paradigm for solving combinatorial optimization problems by shifting from traditional, fixed construction heuristics to dynamically generated, instance-specific policies. We retain the foundational step-by-step framework of construction methods in which a solution is built incrementally. However, we fundamentally revolutionize the core decision-making logic. Instead of relying on a universal, handcrafted rule (e.g., Nearest Neighbor), which is often myopic and susceptible to poor initial choices leading to local optima, our method leverages a LLM to synthesize a sophisticated decision-making function on-the-fly for each problem instance. At every construction step $t$, the LLM acts as a reasoning engine. It is provided with the complete state of the problem, including the current partial tour $S_{t}$, the set of unvisited candidate nodes $U_{t}$, the global distance matrix $C$, and the tour’s origin node to facilitate reasoning about the final closing cost [9, 25].
The advantage of HIFO-Prompt lies in its ability to offload the complex, knowledge-intensive task of heuristic design to the LLM. Through carefully engineered prompts, we guide the model not merely to select a node but to generate a computational policy that embodies advanced strategic principles. The synthesized policy can perform non-trivial reasoning, such as conducting evaluations of multi-step consequences to look beyond immediate greedy gains. It can be guided to implement mechanisms analogous to maintaining a belief over a set of promising candidates, exploring their short-term implications before committing to a single path. Furthermore, the generated function can incorporate stochastic simulations or rollouts to approximate the long-term value of different choices, mimicking the exploration capabilities of advanced search algorithms.
Guided Local Search
Improvement heuristics, such as local search, start with a complete tour and iteratively refine it. Guided Local Search (GLS) is an advanced metaheuristic that enhances local search by introducing a guidance mechanism to escape local optima [72, 66, 62]. GLS achieves this by modifying the objective function with penalties on certain solution features that appear in locally optimal solutions [73]. For the Traveling Salesperson Problem (TSP), the most natural features to penalize are the edges of the tour [74].
The standard GLS cost function is augmented with a penalty term:
$L_{\text{aug}}(s)=L(s)+\lambda\sum_{(u,v)\in s}p_{uv}$ (15)
where $L(s)$ is the original tour length, the sum is over all edges $(u,v)$ in tour $s$, $p_{uv}$ is a penalty counter for using the edge between cities $u$ and $v$, and $\lambda$ is a regularization parameter. An efficient implementation involves creating a penalized distance matrix $D^{\prime}$ for the local search:
$D^{\prime}_{uv}=d_{uv}+\lambda\cdot p_{uv}\quad.$ (16)
Minimizing the tour length using $D^{\prime}$ is equivalent to minimizing $L_{\text{aug}}(s)$. The critical challenge lies in designing the rule for updating the penalty matrix $P=\{p_{uv}\}$. When the local search becomes trapped in a local optimum $s^{*}$, a state-dependent update heuristic, $H_{\text{update}}$, is invoked to determine which edges in $s^{*}$ should be penalized:
$P_{\text{new}}=H_{\text{update}}(P_{\text{old}},s^{*},D,N)\quad,$ (17)
where $N$ is a matrix of edge usage frequencies. Typically, this heuristic is a static, handcrafted rule that increments the penalties for a subset of edges in $s^{*}$. This update reshapes the search landscape to guide the search away from the current basin of attraction.
HiFo-Prompt automates the discovery of the update heuristic $H_{\text{update}}$. Instead of relying on a static, human-designed rule, we task a LLM with synthesizing a complete, executable Python function to serve as $H_{\text{update}}$.
At each GLS iteration, after converging to a local optimum $s^{*}$, this LLM-generated function is invoked. It receives the full state necessary for intelligent penalization—the current penalty matrix ($P_{\text{old}}$), the local optimum tour ($s^{*}$), the original distance matrix ($D$), and the edge frequency matrix (NN)—and outputs a new penalty matrix, $P_{\text{new}}$. This updated matrix then modifies the search costs via Eq. 16 for the next major iteration. The local search itself is driven by fundamental prompt strategies like Relocate [75] and 2-opt [76]. Our contribution lies not in these prompt strategies but in the automated design of the sophisticated $H_{\text{update}}$ function through HiFo-Prompt, which provides the intelligence to guide these prompt strategies effectively.
We further elevate this concept by embedding heuristic generation within an evolutionary framework. We treat the distinct $H_{\text{update}}$ functions as a population of individuals. The LLM itself serves as the primary genetic operator. Through structured prompts for crossover and mutation, the LLM intelligently combines or modifies existing high-performing strategies to produce novel offspring [9].
B.2 Online Bin Packing Problem
The Online Bin Packing Problem (OBP) [48] is a classic sequential decision-making problem. We are presented with a sequence of items, $A=(a_{1},a_{2},\dots,a_{T})$, arriving one at a time. Each item $a_{t}$ has a size $s_{t}\in(0,1]$. We have an unlimited supply of bins, each with a unit capacity of 1. The core constraints of the OBP are:
  • Online Constraint: When item $a_{t}$ arrives, a decision must be made to place it into a bin without any knowledge of future items ($a_{t+1},\dots,a_{T}$).
  • Irrevocable Placement: Once an item is placed in a bin, it cannot be moved.
  • Capacity Constraint: For any bin $B_{j}$, the sum of the sizes of all items placed within it must not exceed 1.
The objective is to minimize the total number of bins used after all $T$ items have been placed [77, 78]. If we let $y_{j}=1$ if bin $j$ is used and $y_{j}=0$ otherwise, the goal is to minimize $\sum_{j}y_{j}$ [79].
Given the online nature of the problem, optimal solutions are generally not achievable. Instead, high-performance algorithms rely on sophisticated greedy placement policies or heuristics. Following the approach of [8], we frame the task as learning a superior placement heuristic. Specifically, we use the HiFo-Prompt framework to design a scoring function, $H_{\text{score}}$, that determines the most suitable bin for an incoming item.
When an item $a_{t}$ with size $s_{t}$ arrives, the HiFo-Prompt-generated heuristic is invoked. It takes as input the item’s size and the state of all currently open bins. The state of a bin $B_{j}$ is captured by its residual capacity, $c_{j}=1-\sum_{a_{k}\in B_{j}}s_{k}$. The scoring function produces a scalar value for each candidate bin:
$\sigma_{j}=H_{\text{score}}(s_{t},c_{j})$ (18)
where $\sigma_{j}$ represents the "desirability" of placing item $a_{t}$ into bin $B_{j}$. The placement policy is then to assign the item to the valid bin with the highest score:
$j^{*}=\underset{j\,|\,c_{j}\geq s_{t}}{\arg\max}\,\sigma_{j}$ (19)
If no existing bin can accommodate the item (i.e., the set of valid bins is empty), a new bin is opened. The intelligence of our method lies in HiFo-Prompt’s ability to discover a non-trivial scoring function $H_{\text{score}}$ that implicitly balances competing objectives, such as leaving space for potentially larger future items versus consolidating small items efficiently. This contrasts with classic heuristics like Best Fit (which is equivalent to $H_{\text{score}}(s_{t},c_{j})=-c_{j}$) or First Fit, by allowing for a much richer and more adaptive decision-making process.
B.3 Flow Shop Scheduling Problem
The Flow Shop Scheduling Problem (FSSP) [49, 80] is a canonical NP-hard scheduling challenge. The task is to schedule a set of $n$ jobs, $J=\{1,\dots,n\}$, on a series of $m$ machines, $M=\{1,\dots,m\}$. Each job $i\in J$ requires $m$ operations, with the $j$-th operation occurring on machine $j$. The processing time for job $i$ on machine $j$ is given by $T_{ij}$ from a processing time matrix $T$ [81, 82]. A solution [83] is a permutation of jobs, $\pi=(\pi_{1},\dots,\pi_{n})$, which dictates the processing order on all machines. Key constraints include that no machine can process multiple jobs at once, and no job can be on multiple machines simultaneously.
The objective is to find a permutation $\pi^{*}$ that minimizes the makespan, $C_{\max}$. This is the total time elapsed until the last job completes its final operation. The completion time $C(\pi_{i},j)$ for job $\pi_{i}$ on machine $j$ is calculated recursively:
$C(\pi_{i},j)=\max\big{(}C(\pi_{i-1},j),C(\pi_{i},j-1)\big{)}+T_{\pi_{i},j}$ (20)
with base cases $C(\pi_{0},j)=0$ and $C(\pi_{i},0)=0$. The makespan for a sequence $\pi$ is therefore $C_{\max}(\pi)=C(\pi_{n},m)$.
Due to the problem’s complexity, we employ a local search metaheuristic [84, 9]. To prevent the search from being trapped in local optima, we use the HiFo-Prompt framework to design a sophisticated guidance strategy automatically. When the search converges to a locally optimal sequence $\pi^{*}$, HiFo-Prompt-generated heuristic, $H_{\text{guide}}$, is invoked. It takes the current sequence, the original processing time matrix, and problem dimensions to produce both a new time matrix $T^{\prime}$ and a designated list of jobs to perturb, $J_{\text{perturb}}$:
$(T^{\prime},J_{\text{perturb}})=H_{\text{guide}}(\pi^{*},T,n,m)$ (21)
This dual output provides a powerful guidance mechanism. The new matrix $T^{\prime}$ reshapes the search landscape by penalizing attributes of the local optimum, effectively steering the search toward unexplored regions. Concurrently, the list $J_{\text{perturb}}$ directs subsequent local search operators, such as insertions or swaps, to focus their computational effort on a specific subset of critical jobs. This combined strategy of altering the problem’s cost structure while focusing on the search operators constitutes a complete and intelligent guidance component, designed automatically by our framework.
B.4 Bayesian Optimization
Bayesian Optimization (BO) [50] and its ongoing development are of paramount importance, as it provides the leading framework for sample-efficiently navigating the complex, high-cost search spaces prevalent in modern science and engineering. Bayesian Optimization (BO) has emerged as a principal framework for this task, excelling in applications like hyperparameter tuning and automated scientific discovery. The power of BO lies in its sample efficiency. It builds a probabilistic surrogate model of the objective function. It then uses an acquisition function [85, 86] to intelligently decide where to sample next, thereby minimizing the number of costly evaluations.
Our work addresses a particularly demanding variant: cost-aware BO [87, 88, 15]. In this setting, each function evaluation $f(\mathbf{x})$ has a heterogeneous and unknown cost, denoted by $c(\mathbf{x})$. The goal is to find the global maximum of $f(\mathbf{x})$ within a fixed total budget $B_{\text{total}}$. This requires sequentially choosing evaluation points $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ to maximize the final outcome, subject to the constraint that $\sum_{i=1}^{N}c(\mathbf{x}_{i})\leq B_{\text{total}}$ [89]. To manage this, the BO agent maintains two surrogate models, typically Gaussian Processes (GPs) [90]. One GP models the objective function, predicting its posterior mean $\mu_{f}(\mathbf{x})$ and standard deviation $\sigma_{f}(\mathbf{x})$. A second GP models the evaluation cost, providing a cost prediction $\mu_{c}(\mathbf{x})$.
The core intelligence of the agent is encoded in its acquisition function, $\alpha(\mathbf{x})$. This function must navigate a complex trade-off between seeking high rewards (exploitation), reducing model uncertainty (exploration), and managing evaluation costs. At each iteration $t$, the next evaluation point $\mathbf{x}_{t+1}$ is selected by maximizing this utility:
$\mathbf{x}_{t+1}=\arg\max_{\mathbf{x}\in\mathcal{X}}\alpha(\mathbf{x}|\mathcal{D}_{t},B_{\text{rem}})$ (22)
where $\mathcal{D}_{t}=\{(\mathbf{x}_{i},y_{i},c_{i})\}_{i=1}^{t}$ is the set of previously evaluated points and $B_{\text{rem}}$ is the remaining budget. The design of $\alpha(\mathbf{x})$ is the single most critical factor for achieving high performance.
We propose to automate the discovery of superior acquisition functions using the HiFo-Prompt framework. Rather than relying on static, human-designed heuristics, HiFo-Prompt generates a novel utility function, $H_{\text{utility}}$, from scratch. This generated function is highly context-aware, synthesizing all critical information available at each decision step. It explicitly considers the surrogate models’ predictions, the best-found solution so far ($y_{t}^{*}=\max_{i}y_{i}$), and the dynamic state of the budget:
$\alpha(\mathbf{x})=H_{\text{utility}}\big{(}\mu_{f}(\mathbf{x}),\sigma_{f}(\mathbf{x}),\mu_{c}(\mathbf{x}),y_{t}^{*},B_{\text{used}},B_{\text{total}}\big{)}$ (23)
By generating a holistic function that reasons about the interplay between potential gain, uncertainty, cost, and remaining resources, HiFo-Prompt creates powerful and adaptive sampling strategies. This approach moves beyond hand-crafted designs, enabling superior performance in complex, budget-constrained optimization scenarios.
Appendix C More Results
C.1 Traveling Salesman Problem
To demonstrate the generality and robustness of the heuristic designed by our proposed method, we conduct a comprehensive evaluation using the real-world benchmark dataset TSPLib [59]. TSPLib is a well-established collection of the TSP instances, widely used in the research community for benchmarking optimization algorithms. For our experiments, we select a diverse subset of TSP instances from TSPLib, specifically focusing on those containing no more than 500 nodes. This selection criterion ensures a manageable problem scale while still retaining the complexity necessary to assess algorithmic performance effectively.
name EoH MCTS ReEvo Hsevo Ours
eil51 7.665% 17.133% 4.409% 4.295% 6.691%
berlin52 16.080% 17.481% 15.678% 19.814% 10.973%
eil76 10.566% 15.193% 8.450% 10.034% 8.439%
pr76 25.222% 29.371% 21.089% 21.718% 20.505%
kroA100 24.347% 24.315% 17.509% 18.758% 16.616%
kroB100 30.274% 22.593% 9.059% 13.266% 20.502%
kroC100 25.542% 10.212% 27.343% 36.176% 10.000%
kroD100 31.497% 25.765% 25.797% 13.258% 32.315%
kroE100 24.618% 21.299% 23.905% 22.218% 14.925%
rd100 12.654% 11.470% 8.660% 12.995% 7.921%
eil101 14.421% 22.947% 12.163% 12.372% 7.229%
lin105 40.297% 16.797% 27.774% 42.519% 13.974%
pr107 9.900% 5.004% 7.894% 7.537% 3.811%
ch130 21.406% 6.930% 7.352% 13.107% 7.090%
ch150 19.604% 9.874% 12.644% 10.185% 4.275%
kroA150 27.891% 20.517% 25.887% 21.869% 20.070%
kroB150 22.403% 27.716% 29.499% 32.729% 13.677%
pr152 12.302% 10.457% 12.473% 11.960% 6.545%
u159 13.444% 7.074% 7.951% 13.644% 7.828%
kroA200 26.395% 26.219% 27.366% 28.537% 22.702%
kroB200 20.980% 20.083% 23.861% 21.590% 14.314%
tsp225 32.938% 20.193% 21.703% 22.503% 18.683%
a280 34.081% 24.095% 27.921% 31.736% 15.936%
rd400 14.612% 14.745% 13.865% 15.213% 8.075%
d493 18.833% 13.248% 14.776% 21.179% 7.976%
mean 21.519% 17.629% 17.401% 19.168% 12.843%
Table 6: Results on TSPLib. The best results are highlighted in bold.
In our evaluation framework, we adopt a step-by-step construction approach to solve the TSP, which incrementally builds a tour by selecting the next node based on a learned heuristic. This paradigm allows us to evaluate the quality of decisions made at each step and better observe the contribution of the designed heuristic to the final solution quality. To rigorously assess the effectiveness and competitiveness of our proposed heuristic, we compare its performance against state-of-the-art LLM-based AHD baselines. These methods represent recent advances in leveraging large language models for combinatorial optimization tasks and serve as strong comparative baselines in our study. The experimental results, which include performance metrics, are summarized in Table 6. These results provide empirical evidence that our method not only performs competitively but also generalizes well across a variety of TSP instances in real-world scenarios.
C.2 Online Bin Packing Problem
To provide a more comprehensive empirical validation and to assess the robustness of our framework, we conducted further evaluations of the online Bin Packing Problem (BPP) on a broader and more challenging set of Weibull-distributed instances. These instances, which more closely mimic real-world scenarios than uniform distributions, were generated following the protocol established in prior work on LLM-based heuristic discovery  [8]. Table 7 presents a detailed comparative analysis of our method against a wide spectrum of competitors, including both classical, widely-adopted handcrafted heuristics (First Fit, Best Fit) and a suite of contemporary approaches based on Large Language Models.
For each combination of bin capacity and problem size, the dataset includes five unique instances. Performance is quantified by the average percentage gap to the known theoretical lower bound across these instances, where a smaller value signifies a more efficient and effective packing solution. The empirical results unequivocally demonstrate the superior performance of the heuristic we have discovered. As shown in the table, our method consistently outperforms all baseline methods in nearly every setting, securing the smallest average gap in 8 out of the 9 distinct configurations tested. This highlights a remarkable level of consistency and dominance. The only exception is the (Capacity=100, Size=10k) case, where Funsearch achieves a marginally lower gap. However, our method’s strong performance across the entire parameter space underscores its greater reliability and generalizability compared to methods that may excel only in specific, narrow scenarios. Notably, our approach scales gracefully, achieving extremely low gap values (e.g., 0.41%, 0.42%) on the largest and most complex problems, significantly surpassing both traditional algorithms and other state-of-the-art LLM-driven frameworks. This comprehensive evaluation on challenging instances further validates the efficacy of our evolutionary framework, showcasing its capacity to discover sophisticated and high-performance heuristics that are robust across varied problem characteristics.
Capacity Size First Fit* Best Fit* Funsearch* EoH ReEvo MCTS-AHD HSEvo Ours
100 1k 5.32% 4.87% 3.78% 4.43% 3.63% 3.38% 3.53% 1.99%
5k 4.40% 4.08% 0.80% 1.07% 0.78% 0.99% 0.79% 0.66%
10k 4.44% 4.09% 0.33% 0.71% 0.35% 0.84% 0.34% 0.44%
300 1k 4.93% 4.48% 4.93% 4.48% 7.34% 3.21% 8.03% 2.01%
5k 4.18% 3.83% 1.07% 1.07% 4.47% 0.95% 5.41% 0.63%
10k 4.20% 3.87% 0.49% 0.73% 4.05% 0.85% 5.01% 0.41%
500 1k 4.97% 4.50% 6.75% 4.50% 5.92% 3.20% 6.45% 1.95%
5k 4.27% 3.91% 1.47% 1.06% 3.24% 0.95% 4.04% 0.66%
10k 4.28% 3.95% 0.74% 0.70% 2.79% 0.85% 3.49% 0.42%
Table 7: Results on Online BPP in Weibull instances with varied capacities and problem sizes. Results marked with * denote values taken from  [9].
C.3 Flow Shop Scheduling Problem
Table 8 presents a comparative analysis of the performance of EoH and our proposed method on FSSP of varying scales, defined by the number of jobs $n$ and machines $m$. Performance is evaluated using two key metrics: the objective function value and the relative gap, where the gap is measured with respect to the solution quality of an advanced handcrafted heuristic. A smaller gap reflects a solution that is closer to the heuristic baseline and, therefore, indicates higher solution quality. The results demonstrate that our method consistently achieves smaller gaps across all tested problem configurations, regardless of scale. Moreover, in many cases, it outperforms the advanced heuristic itself in terms of the raw objective value. This consistent superiority suggests that our approach not only generalizes well across diverse FSSP instances [9] but also offers a competitive alternative to domain-specific heuristics, exhibiting both strong effectiveness and robustness.
$n$ $m$ EoH Ours
Gap Obj. Gap Obj.
20 5 0.21% 1225.1 -0.01% 1222.4
10 0.24% 1517.4 0.15% 1515.9
20 0.18% 2239.1 0.10% 2237.4
50 5 0.02% 2736.8 0.00% 2736.4
10 0.29% 3006.2 0.12% 3001.2
20 0.75% 3759.1 0.60% 3753.4
100 5 -0.04% 5244.8 -0.05% 5244.5
10 0.17% 5639.8 0.09% 5635.3
20 0.93% 6421.6 0.47% 6392.6
200 10 0.25% 10699.3 0.08% 10681
20 0.94% 11480.7 0.63% 11444.9
Table 8: Results on FSSP. The best results are highlighted in bold.
C.4 Bayesian Optimization
Table 9 presents the experimental results on the design of cost-aware acquisition functions (CAFs) [88] in Bayesian Optimization. We compare our method against both manually crafted CAFs and those generated by LLM-based AHD methods. To ensure a fair comparison, all LLM-based AHD approaches utilize the same underlying language model. Our method achieves superior performance on the majority of benchmark functions, outperforming traditional CAFs (e.g., EI, EIpu, EI-cool) [88] as well as recent LLM-based AHD baselines (e.g., EoH [9], MCTS [15]). The performance advantage is particularly pronounced on challenging functions such as Rastrigin, ThreeHumpCamel, and Shekel. These results underscore the robustness and strong generalization ability of our approach across a wide range of optimization landscapes.
Ackley Rastrigin Griewank Rosenbrock Levy ThreeHumpCamel
EI* 2.66 4.74 0.49 1.26 0.01 0.05
EIpu* 2.33 5.62 0.34 2.36 0.01 0.12
EI-cool* 2.74 5.78 0.34 2.29 0.01 0.07
EOH 3.11 3.48 0.72 2.57 0.04 0.18
MCTS 3.23 0.87 0.43 1.30 0.01 0.05
Ours 1.78 0.45 0.41 1.50 0.00 0.01
StyblinskiTang Hartmann Powell Shekel Hartmann Cosine8
EI* 0.03 0.00 18.89 7.91 0.03 0.47
EIpu* 0.02 0.00 19.83 7.92 0.03 0.47
EI-cool* 0.03 0.00 14.95 8.21 0.03 0.54
EOH 2.89 0.01 13.71 8.71 0.47 1.04
MCTS 0.02 0.00 1.91 5.14 0.08 0.29
Ours 0.02 0.01 2.65 4.08 0.57 0.28
Table 9: Results on BO. The results denote the absolute error relative to the optimal solution. Results marked with * are from  [88]. The best results are highlighted in bold.
Appendix D Algorithm Details
This section provides detailed pseudocode for the core components of the HiFo-Prompt framework. We first present the main algorithm, followed by the specific implementations for the survival selection and parent selection mechanisms.
D.1 Main Framework Algorithm
Algorithm 1 outlines the complete, end-to-end procedure for the HiFo-Prompt framework. It details the main evolutionary loop, starting from population initialization and proceeding through iterative generations. Each generation consists of four primary phases: (1) Foresight and Hindsight, where the Evolutionary Navigator and Insight Pool generate the control vector; (2) Construction, where the LLM generates new heuristics under the guidance of the control vector; (3) Evaluation & Feedback, where offspring are evaluated and the Insight Pool learns from the results; and (4) Selection, where the next generation’s population is formed.
Algorithm 1 HiFo-Prompt
Input: Problem definition $\mathcal{F}$, optional seed algorithms $\mathcal{P}_{\text{seed}}$
Parameter: Population size $N_{\text{pop}}$, number of generations $G_{\max}$, set of evolutionary foundation prompt strategies $\mathcal{O}$ with weights $\mathcal{W}$, number of parents $m$
Output: The final population of evolved algorithms $\mathcal{P}$
1: Initialize Population $\mathcal{P}$:
2:if $\mathcal{P}_{\text{seed}}$ is provided then
3:  $\mathcal{P}\leftarrow$ Evaluate and populate from $\mathcal{P}_{\text{seed}}$
4:else
5:  $\mathcal{P}\leftarrow\emptyset$
6:  while $|\mathcal{P}|<N_{\text{pop}}$ do
7:   $\text{Ind}_{\text{new}}\leftarrow$ Generate an initial algorithm via LLM
8:   Evaluate fitness of $\text{Ind}_{\text{new}}$ using $\mathcal{F}$
9:   Add valid $\text{Ind}_{\text{new}}$ to $\mathcal{P}$
10:  end while
11:end if
12: Initialize Insight Pool $\mathcal{M}_{\text{insight}}$ with a set of default design principles
13: Initialize Evolutionary Navigator $\mathcal{C}_{\text{meta}}$
14:
15:for $g=1$ to $G_{\max}$ do
16:  Update $\mathcal{C}_{\text{meta}}$ with current population state
17:  if a condition for wisdom extraction is met then
18:   $\mathcal{P}_{\text{top}}\leftarrow$ Select top-performing individuals from $\mathcal{P}$
19:   $\text{new\_tips}\leftarrow$ Prompt LLM to extract generic design principles from $\mathcal{P}_{\text{top}}$
20:   Add new_tips to $\mathcal{M}_{\text{Insight}}$
21:  end if
22:
23:  for each foundation prompt strategy $o\in\mathcal{O}$ do
24:   if $\text{random}()<\mathcal{W}[o]$ then
[…]
26:    $\text{insights}\leftarrow\mathcal{M}_{\text{Insight}}$.get_insights() {Get inspiring principles}
27:    $\mathcal{P}_{\text{parents}}\leftarrow$ Select $m$ parents from $\mathcal{P}$
28:    $\text{Ind}_{\text{new}}\leftarrow$ Generate offspring via LLM using $o$, $\mathcal{P}_{\text{parents}}$, guidance, and insights
29:    Evaluate fitness of $\text{Ind}_{\text{new}}$ using $\mathcal{F}$
30:    if $\text{Ind}_{\text{new}}$ is valid and not duplicated then
31:     $\text{effectiveness}\leftarrow$ Calculate effectiveness of tips based on $\text{Ind}_{\text{new}}$’s performance
32:     $\mathcal{M}_{\text{Insight}}$.update_insight_stats(insights, effectiveness) {Feedback loop}
33:     Add $\text{Ind}_{\text{new}}$ to $\mathcal{P}$
34:    end if
35:   end if
36:  end for
37:  $\mathcal{P}\leftarrow$ SurvivalSelection($\mathcal{P}$, $N_{\text{pop}}$) {Manage population size}
38:end for
39:return Best individuals from $\mathcal{P}$
D.2 Survival Selection Mechanism
To maintain a constant population size across generations, we employ a survival selection mechanism after new offspring have been generated and evaluated. Algorithm 2 details this procedure. It implements a greedy, elitist strategy that first filters out any invalid or unsuccessfully evaluated individuals. It then ensures uniqueness based on objective values to maintain diversity in the performance space. Finally, it sorts the unique, valid individuals by their performance and truncates the population to the target size $N_{target}$, ensuring that only the highest-performing heuristics are retained for the next generation.
Algorithm 2 Greedy Population Management
Input: A population of individuals $P$, Target population size $N_{target}$
Output: A new population $P_{new}$ with size at most $N_{target}$
[…]
2: Initialize an empty list $P_{unique}$
3: Initialize an empty set of seen objectives $O_{seen}$
4:for each individual $i\in P_{valid}$ do
5:  if $i.objective\notin O_{seen}$ then
6:   Add $i$ to $P_{unique}$
7:   Add $i.objective$ to $O_{seen}$
8:  end if
9:end for
[…]
11: Sort $P_{unique}$ in ascending order based on objective values.
[…]
13:return $P_{new}$
D.3 Parent Selection Mechanism
For the generative operators that require parent heuristics (e.g., crossover and mutation), a parent selection mechanism is used to choose individuals from the current population. Algorithm 3 describes the rank-based selection method used in our framework. This approach assigns a selection probability to each individual that is inversely proportional to its performance rank. This ensures that higher-performing individuals are more likely to be selected as parents, creating a strong selection pressure that guides the search towards promising regions of the heuristic space, while still allowing lower-ranked individuals a chance to contribute.
Algorithm 3 Rank-Based Parent Selection
Input: Sorted population $P$ (best to worst), Number of parents to select $m$
Output: A list of selected parents $P_{parents}$
1: Let $N\leftarrow$ size of $P$
2: Initialize an empty list for probabilities $W\leftarrow\{\}$
3:for $k=0$ to N1N-1 do
4:  Calculate probability $w_{k}\propto 1/(k+1+N)$ {Rank-based weighting}
5:  Append $w_{k}$ to $W$
6:end for
7: Select $m$ individuals from $P$ with replacement, using probabilities $W$, to form $P_{parents}$.
8:return $P_{parents}$
Appendix E Limitation and Future Work
E.1 Limitation
While HiFo-Prompt demonstrates significant advancements in automated heuristic design, its current architecture possesses inherent limitations that define the boundaries of its present capabilities.
First, the core decision-making mechanism of the Evolutionary Navigator is predicated on a static, handcrafted control logic. This rule-based system, while interpretable, lacks the capacity for self-adaptation. Its fixed thresholds for stagnation, progress, and diversity are calibrated for the evaluated problems but may not be universally optimal, potentially constraining its performance on novel problem landscapes or over different evolutionary timescales. The Navigator can react to predefined states but cannot learn or refine its control strategy from experience.
Second, the knowledge evolution within the Insight Pool is fundamentally intra-task. The framework excels at capturing, refining, and reusing design principles within the context of a single optimization problem. However, the true generalizability of these learned insights across different problem domains remains an unevaluated and open question. We have not yet systematically validated whether insights distilled from solving one problem class can effectively bootstrap the learning process on a structurally different, unseen one.
E.2 Future Work
The limitations mentioned above naturally chart a course for several high-impact avenues for future research, aimed at enhancing the framework’s autonomy, generality, and strategic depth.
A primary research thrust will be to transcend the Evolutionary Navigator’s current heuristic-driven design by developing it into a learned metacontroller. We propose parameterizing its control function and leveraging techniques from the domain of Meta-Reinforcement Learning (Meta-RL). In this paradigm, the Navigator would operate as a high-level agent, learning a control policy that dynamically maps observational data from the evolutionary process—such as population diversity metrics, fitness landscape topology, and rates of convergence—to a nuanced control vector. This vector would modulate critical evolutionary parameters in real-time, including operator probabilities, selection pressure, and even strategic alterations to the LLM prompts themselves. The reward signal for this meta-agent would be carefully engineered to reflect overarching goals of evolutionary efficiency, such as maximizing the rate of fitness improvement or minimizing the computational cost to reach a performance threshold. Successfully implementing this would elevate the framework from a system guided by static rules to one capable of learning and deploying its own adaptive control strategies online, thereby achieving a superior order of autonomy and problem-specificity.
Another critical, complementary direction is the systematic investigation of inter-task knowledge transfer, with the goal of evolving the framework from a single-task solver into a more general algorithmic discovery platform. This research will proceed along two parallel vectors. First, we will conduct a rigorous empirical evaluation of the framework’s zero-shot and few-shot transfer capabilities. By applying a mature Insight Pool—developed for a source task—to novel target domains, we can precisely quantify the generality of the learned principles and measure the extent to which discovery costs can be amortized across problems. Second, we will pioneer the exploration of more abstract and powerful knowledge representations that move beyond the inherent ambiguities of a natural language string. This includes investigating structured canonical forms, such as predicate logic or probabilistic program sketches, and rich semantic representations like knowledge graphs. The central hypothesis is that such formalisms can more effectively decouple a core algorithmic invariant from its domain-specific instantiation, thereby creating a more robust foundation for seamless and compositional cross-domain knowledge transfer.
Appendix F Prompts with Foresight and Hindsight
This section provides the specific, concrete templates of the prompts used to guide the LLM, offering a transparent and reproducible view of the core interaction engine at the heart of our HiFo-Prompt framework. The ’HiFo’ (Hindsight-Foresight) name is not arbitrary; it directly reflects our core methodology: the strategic deployment of distinct, context-aware prompts at different stages of the evolutionary process. ’Foresight’ prompts are strategically employed during the initial generation and creative mutation stages, encouraging the LLM to explore a diverse space of novel heuristic possibilities. In stark contrast, ’Hindsight’ prompts play a critical role in reflection and data-driven refinement, leveraging empirical performance feedback from past evaluations to methodically prune the search space and systematically improve promising solutions.
To ground these abstract concepts and make our prompts tangible, we anchor our examples in the canonical Online Bin Packing (OBP) problem. The objective in OBP is to assign an incoming sequence of items of varying sizes into a minimum number of fixed-capacity bins, with the crucial constraint that each item must be placed without knowledge of future items. Within this problem context, our framework’s specific task is to evolve a high-performance Python scoring function, score(item, bin), which acts as the core decision-making logic. This function evaluates the suitability of a specific candidate bin for an incoming item. A higher returned score signifies a more desirable placement, and the overarching control logic of the framework places the item in the bin that yields the maximum score. The following subsections provide verbatim templates for each distinct phase of the evolutionary process, clearly demonstrating how the system’s guidance dynamically shifts its focus—from broad exploration to focused exploitation—as the search progresses.
F.1 Initial Prompt Strategy I1
The i1 operator uses the following prompt template to generate the initial population of heuristics. The guidance provided to the LLM at this stage comes from the initial state of the framework’s components, including a set of high-quality seed insights.
Prompt for Operator i1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. First, describe your new algorithm and main steps in one sentence. The description must be inside a brace. Next, implement it in Python as a function named score. This function should accept 2 input(s): ’item’, ’bins’. The function should return 1 output(s): ’scores’. The score function is designed to evaluate the placement options for a given item. It takes the item to be placed and the current list of bins as input. It returns a list of numerical scores, with each score corresponding to a bin in the input list. This list of scores guides the heuristic in selecting the most suitable bin for the item according to the generated logic. Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle...> For the evolutionary regime, please pay special attention to: <A specific instruction from Design Directive> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
F.2 Recombination Prompt Strategy E1
The following template for the e1 operator is a representative example of a prompt used during the main evolutionary loop. It demonstrates how parent heuristics and the full, dynamic guidance from the meta-cognitive components are integrated.
Prompt for Operator e1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have k existing algorithms with their codes as follows: No.1 algorithm and the corresponding code are:
<Description of the first algorithm>
<The Python code implementation of the first algorithm>
No.k algorithm and the corresponding code are:
<Description of the last algorithm>
<The Python code implementation of the last algorithm>
Please help me create a new algorithm that has a totally different form from the given ones. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle...> For the evolutionary regime, please pay special attention to: <A specific Design Directive for promoting structural novelty, preserving diversity, and avoiding premature convergence> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
F.3 Recombination Prompt Strategy E2
The e2 operator focuses on "motivated recombination." It prompts the LLM to first identify a common "backbone" or core principle shared by the parent heuristics and then to create a new, improved algorithm based on that shared foundation.
Prompt for Operator e2 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have k existing algorithms with their codes as follows: No.1 algorithm and the corresponding code are:
<Description of the first algorithm>
<The Python code implementation of the first algorithm>
No.k algorithm and the corresponding code are:
<Description of the last algorithm>
<The Python code implementation of the last algorithm>
Please help me create a new algorithm that has a totally different form from the given ones but can be motivated from them. Firstly, identify the common backbone idea in the provided algorithms. Secondly, based on the backbone idea, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Thirdly, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> For this recombination, please pay special attention to: <A specific Design Directive for simplification prune low-impact features> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
F.4 Mutation Prompt Strategy M1
The M1 operator implements a form of targeted mutation, generating a single-parent descendant by refining its most critical components. It performs surgical modifications to elements like scoring rules or parameter weights, while meticulously safeguarding the parent’s established, high-performing logic. This process is not random; it is guided by a synthesis of recent, high-utility design "insights" and a high-level strategic directive. By injecting controlled, purposeful diversity, M1 enables the search to escape local optima without dismantling the parent’s effective architecture. This deliberate balance between exploitation (refining what works) and exploration (seeking novelty) is crucial for accelerating convergence towards superior heuristics.
F.5 Mutation Prompt Strategy M2
The M2 operator implements parameter mutation, a targeted process to modify a heuristic’s numerical behavior. It instructs the LLM to first deconstruct the parent’s score function to identify its key hyperparameters. Following this analysis, the model is prompted to generate a new set of parameter values. This generation can either explore novel configurations through significant changes or fine-tune existing ones via subtle adjustments. This surgical approach methodically injects parametric diversity into the population while preserving the integrity of the core algorithmic logic.
F.6 Mutation Prompt Strategy M3
The M3 operator is designed for structural simplification to enhance heuristic robustness and combat overfitting. It instructs the LLM to perform a critical analysis of the parent score function, specifically targeting components suspected of being over-specialized to in-distribution data. These potentially brittle or overly complex segments are then strategically pruned or streamlined. The outcome is a more parsimonious and computationally lean implementation that is theorized to exhibit superior generalization to out-of-distribution scenarios, all while preserving the original function signature to ensure architectural compatibility.
Prompt for Operator m1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Algorithm description: <Description of the parent algorithm> Code: <The Python code implementation of the parent algorithm> Please assist me in creating a new algorithm that has a different form but can be a modified version of the algorithm provided. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> For this mutation, please pay special attention to: <A specific Design Directive for mutation> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
Prompt for Operator m2 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Algorithm description: <Description of the parent algorithm> Code: <The Python code implementation of the parent algorithm> Please identify the main algorithm parameters and assist me in creating a new algorithm that has different parameter settings of the score function provided. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> When adjusting parameters, please pay special attention to: <A specific Design Directive for parameter tuning> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
Appendix G Managing Foresight and Hindsight Knowledge
G.1 Hindsight Evolution via Insight Distillation
To continually enrich our system’s understanding of effective optimization strategies, we employ a Large Language Model (LLM) to distill high-level design principles from high-performing algorithms. This process is guided by a structured prompt, shown below, which provides the LLM with the descriptions and/or code of elite solutions discovered during an evolutionary run. The core instruction tasks the model with synthesizing concise, generalizable, and performance-positive patterns, drawing insights from both the conceptual descriptions and the code implementations. This automated extraction mechanism allows our system to learn from its own successes and progressively build a more sophisticated knowledge base.
Prompt for Operator m3 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Code: <The Python code implementation of the parent algorithm> First, identify the main components in the function above. Next, analyze which of these components may be overfitting to the in-distribution instances. Then, simplify those components to enhance generalization to out-of-distribution cases. Finally, provide the revised code, keeping the function name, inputs, and outputs unchanged. Provide the complete revised function implementation, preserving its original signature. <The function name, inputs & outputs specification from your prompt> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> When simplifying, please pay special attention to: <A specific Design Directive for simplification> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
Prompt Template for Insight Extraction The following are core descriptions and/or code of high-performance optimization algorithms evolved recently: Algorithm 1: <Natural language description and/or code of elite individual 1> Algorithm 2: <Natural language description and/or code of elite individual 2> Algorithm n: ... (and so on for the top 30% of the population) Please extract 1-2 concise, generic, and performance-positive [design principles] or [effective patterns] from the above algorithms. These principles should be applicable to various combinatorial optimization problems, not just the specific problem domain. When formulating these principles, it is essential to draw insights from both the conceptual natural language descriptions and their corresponding code implementations. Focus on identifying the underlying strategic design choices and algorithmic methodologies rather than superficial characteristics or specific implementation minutiae. Each principle/pattern must be expressed as an independent sentence in the following format: Balance local optimization with global solution structure when making decisions. Prioritize choices that maintain flexibility for future decision-making steps. Implement adaptive mechanisms that respond to problem instance characteristics. Provide only the list of principles, without any preamble or other explanatory text.
G.2 Insight Seed Pool
At the inception of the framework’s execution, the LLM bootstraps the heuristic generation process by drawing from a curated repository of Seed Insights, a mechanism designed to mitigate the classic "cold start" problem and channel the model’s creativity. These insights, which encapsulate established principles and canonical rules-of-thumb distilled from decades of human expertise in heuristic design—such as the "Shortest Processing Time" principle in scheduling or the "Nearest Neighbor" concept in routing—are not rigid constraints but rather high-level conceptual guidelines. They are strategically injected into the foundational prompts, often framed within a dedicated "human knowledge" block, to act as an intellectual scaffold. This initial infusion of well-vetted knowledge serves to ground the search, preventing the generation of naive or logically flawed heuristics and providing a potent directional bias that immediately steers the early stages of algorithmic evolution away from vast, unproductive regions of the design space. By ensuring the process begins not from a tabula rasa but from a high-quality, well-founded starting point, these seed insights exert a persistent influence that accelerates convergence and significantly enhances the quality and novelty of all subsequent automated discovery.
Seed Insights Design adaptive hybrid meta-heuristics synergistically fusing multiple search paradigms and dynamically tune operator parameters based on search stage or problem features. Employ machine learning to mine problem structures and use learned insights to intelligently bias towards promising search regions. Explore objective function engineering by introducing auxiliary objectives or dynamically adjusting weights to reshape the search landscape. Construct problem-specialized solution representations and co-design dedicated operators to fully leverage the representation’s structure. Implement intelligent diversification based on solution feature space analysis to systematically target uncovered regions and escape local optima.
G.3 The Directive Pool for Foresight
To operationalize the Evolutionary Navigator’s high-level strategy, we map the chosen regime—Exploration, Exploitation, or Balance—to a specific Design Directive. Each directive is a fine-grained textual instruction, uniformly sampled from a predefined pool corresponding to the active regime. This sampled directive is then integrated into the generation prompt. This two-tiered mechanism facilitates fine-grained control over the generation process, ensuring model outputs are precisely aligned with the overarching evolutionary objective.
Design Directive Balance: Optimizing objective function evaluation criteria. Considering the long-term impact of current decisions. Balancing local optimality with global search strategies. Improving algorithm robustness across different problem instances. Managing computational complexity and time efficiency. Exploitation: Refining core evaluation and scoring functions. Fine-tuning critical algorithm parameters and thresholds. Improving the precision of existing heuristics and rules. Reducing unnecessary computational overhead. Exploration: Exploring novel solution construction methodologies. Investigating alternative problem decomposition approaches. Introducing new randomization or adaptive mechanisms. Experimenting with hybrid strategy combinations.